4x^2-y^2=0
how do you do this hyperbola problem

Found 2 solutions by lwsshak3, richwmiller:
Answer by lwsshak3(11628)

4x^2-y^2=0
how do you do this hyperbola problem
y^2=4x^2
y=±2x
Graphically, given hyperbola degenerates into 2 straight lines with slopes of 2 and -2 that intersect at the origin.

Answer by richwmiller(17219)

that factors into

(2x-y)(2 x+y) = 0
which are two lines that meet